Consumption of Fixed Capital Government Purchases US imports Personal Taxes Transfer Payments US Exports Personal ConsumptionMultiple Choice $940 billion. $975 billion. $804 billion. $1,019 billion.
Answer #1

Answer

GDP(Gross Domestic product) using expenditure method is given by :

GDP = Consumption + gross private domestic investment + Government purchase + Exports - Imports

= 475 + 300 + 315 + 249 - 260

= 1079

(Note : All numbers used till now and further are in billions)

National Income = GDP + Net foreign factor income - Consumption of fixed capital - Statistical discrepancy

= 1079 + 5 - 25 - 40

= 1019 (billion)

Hence, the correct answer is (d) 1019 billion